OpenStudy (anonymous):

Write the first four terms of the geometric sequence in which a1 = -3 and r = -4. a. -3, 12, -48, 192 c. –3, -12, -48, -192 b. 3, 12, 48, 192 d. 3, -12, 48, -192

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (here_to_help15):

@Hero

OpenStudy (here_to_help15):

What grade r u in : P?

OpenStudy (anonymous):

later pls

OpenStudy (phi):

geometric sequence that means the sequence goes \[ a_1 r^0, a_1 r^1, a_2 r^2, a_1 r^3 ,...\]

OpenStudy (anonymous):

@D3xt3R help once again :P

OpenStudy (anonymous):

yeah @phi agree

OpenStudy (phi):

you know a1 and you know r, (and you should know that r^0 (not matter what r is) is 1 so you know the first number in the sequence. what is it ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

oh i see so it is A?

OpenStudy (phi):

in other words, a1*1 = -3*1 = -3 so we can rule out c and d it's 50-50 now!

OpenStudy (phi):

the next term is the first term times r r is -4 what is the first term times -4 ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

A ?

OpenStudy (phi):

only a is on the right track. bingo!

OpenStudy (anonymous):

thanks